# Coresets from Trajectories: Selecting Data via Correlation of Loss Differences

### v1

@@ E.3.1 Samples with Low Loss Curvature (SloCurv)
SloCurv (Garg & Roy, 2023) scores each training sample by an input-loss curvature proxy computed at the end of (proxy) training. For a sample $\vec{z}_{m}$, with model parameters $\theta^{T}$ and random Rademacher directions $v_{r}$ scaled by $h$, the score is
@@ E.3.2 Temporal Dual-Depth Scoring (TDDS)
TDDS (Zhang et al., 2024) builds a coreset by combining two temporal depths of signal from training with a proxy model. Depth 1 computes, for each epoch, the projection of each sample’s per-sample gradient onto the epoch’s accumulated gradient direction. Depth 2 then aggregates these per-epoch contributions over a sliding window of length $J$ and emphasizes their temporal variability (e.g., windowed variance). We maintain windowed statistics in a streaming manner (constant-time updates), so full trajectories need not be stored.
@@ E.3.3 Using Prediction Uncertainty with a Proxy (Dyn-Unc, DUAL)
Dyn-Unc (He et al., 2024) measures prediction uncertainty via a sliding window of length $J$ over per-example target-class probabilities and averages the windowed uncertainty across proxy training.
DUAL (Cho et al., 2025) combines uncertainty with difficulty (window-mean prediction) and computes scores from an early stage of proxy training.
@@ Appendix F Comparison of CLD with Influence
The impact measured by $\mathtt{CLD}$ closely aligns with the “influence" of individual training samples on a model’s predictions that are measured by Training Data Attribution (TDA) methods. TDA methods have been widely employed for tasks such as debugging datasets, interpreting models, and optimizing training efficiency Koh & Liang (2017); Yeh et al. (2018); Feldman & Zhang (2020).
The earliest TDA methods utilized Leave-One-Out (LOO) training, which involves retraining the model after removing specific data points and observing the changes in performance. While straightforward, LOO retraining is computationally prohibitive for modern deep learning models due to the need for multiple retraining cycles Koh & Liang (2017). Recent TDA metrics, such as FZ-Influence ($\mathtt{Infl}$Feldman & Zhang (2020) and $\mathtt{Datamodels}$ Ilyas et al. (2022), have gained popularity owing to precomputed scores for widely-used datasets in computer vision. These methods, however, face scalability challenges.
A prominent alternative that arose was Influence Functions, which estimated the effect of downweighting individual samples using first-order (gradient) and second-order (Hessian) computations Koh & Liang (2017); Basu et al. (2021) performed at the end of training. Methods like $\mathtt{RandSelect}$ Wojnowicz et al. (2016) and $\mathtt{Arnoldi}$ iterations Schioppa et al. (2022) improved computational efficiency by approximating the Hessian. Similarly, $\mathtt{TRAK}$ Park et al. (2023) combined random projections, gradient-based methods, and ensembling to estimate the influence of training samples. However, these approaches often rely on strong assumptions, such as convergence to a unique optimal solution, which limits their applicability to neural networks. Additionally, Hessian computations introduce significant computational overhead. To address these challenges, unrolling-based methods that observe the learning process across training iterations have been proposed. These techniques approximate the impact of samples by differentiating through the optimization trajectory Hara et al. (2019). Among these, $\mathtt{TracIn}$ Pruthi et al. (2020) is a highly efficient method that estimates influence using gradients tracked throughout training. Its practical implementation, $\mathtt{TracInCP}$, uses intermediate checkpoints to alleviate computational burdens. While effective, unrolling methods require storing intermediate training states, leading to high storage and computational costs.
@@
We will now compare the impact measured by this metric ($\mathtt{CLD}_{\texttt{infl}}$) to the influence measured by TDA metrics, by utilizing the linear datamodeling score ($\mathtt{LDS}$) introduced by Park et al. (2023). $\mathtt{LDS}$ measures the correlation between group-level attribution scores ($\mathtt{CLD}_{\texttt{infl}}$ or influence) and their observed impact on model predictions when subsets of training data are used.
$\mathbf{LDS}$ definition For a query data point $z_{q}$, random subsets $\{\mathcal{S}_{j}\}_{j=1}^{C}$ are sampled from the training dataset, where each subset $\mathcal{S}_{j}$ contains $\lceil\alpha N\rceil$ points, with $\alpha\in(0,1)$ as the sampling ratio. Each subset $\mathcal{S}_{j}$ is used to retrain the model $R$ times with different initializations $\{\xi_{r}\}_{r=1}^{R}$ and training parameters $\lambda$, resulting in the model $\theta^{T}_{S_{j},\xi_{r}}$. This trained model is then used to compute a measurable quantity $f(\vec{z}_{q},\theta^{T}_{\mathcal{S}_{j},\xi_{r}})$. A group attribution score, $g_{\tau}(\vec{z}_{q},\mathcal{S}_{j},\mathcal{S})$, is calculated as $g_{\tau}(\vec{z}_{q},\mathcal{S}_{j},\mathcal{S})\coloneqq\sum_{\vec{z}\in\mathcal{S}_{j}}\tau(\vec{z}_{q},\vec{z},\mathcal{S}),$ where $\tau(\vec{z}_{q},\vec{z},\mathcal{S})$ is the attribution score for a training point $\vec{z}$ with respect to $\vec{z}_{q}$. The $\mathtt{LDS}$ is then obtained using Spearman’s rank Spearman (1904) correlation ($\rho_{s}$):
@@
Experimental Setup: We compared the $\mathtt{LDS}$ scores of $\mathtt{CLD}_{\texttt{infl}}$ against those of $\mathtt{TRAK}$, $\mathtt{Arnoldi}$, $\mathtt{TracIn}$, $\mathtt{Infl}$, and $\mathtt{Datamodels}$. Precomputed scores for $\mathtt{Infl}$ and $\mathtt{Datamodels}$ were used for the CIFAR-10 dataset Krizhevsky et al. (2009) with ResNet-9 He et al. (2016), while 10 models were trained for $\mathtt{TRAK}$, $\mathtt{Arnoldi}$, $\mathtt{TracIn}$, and $\mathtt{CLD}_{\texttt{infl}}$. The evaluation employed $C=100$ random subsets, sampling ratios $\alpha$ ranging from $0.3$ to $\frac{N-1}{N}$, a query set of 200 samples, and $R=10$ seeds. The measurable quantity was the accuracy of query samples.
@@
Results and Observations: The results presented in Figure˜8(a) reveal that while the impact captured by $\mathtt{CLD}_{\texttt{infl}}$ is distinct from the influence measured by traditional TDA metrics, it aligns closely with methods such as $\mathtt{TracIn}$ and $\mathtt{TRAK}$ in terms of behavior while being resource-efficient. Notably, the performance gap between these computationally intensive methods and $\mathtt{CLD}_{\texttt{infl}}$ narrows as $\alpha$ increases. The drop in $\mathtt{LDS}$ scores at $\alpha=\frac{N-1}{N}$ is due to the stochastic nature of model retraining [^1] This observation is consistent with the findings of previous research Karthikeyan & Søgaard (2021); Bae et al. (2024)..

### v2

@@ E.3.1 Samples with Low Loss Curvature (SloCurv)
SloCurv (slocurves_garg2023) scores each training sample by an input-loss curvature proxy computed at the end of (proxy) training. For a sample $\vec{z}_{m}$, with model parameters $\theta^{T}$ and random Rademacher directions $v_{r}$ scaled by $h$, the score is
@@ E.3.2 Temporal Dual-Depth Scoring (TDDS)
TDDS (zhang2024_tdds) builds a coreset by combining two temporal depths of signal from training with a proxy model. Depth 1 computes, for each epoch, the projection of each sample’s per-sample gradient onto the epoch’s accumulated gradient direction. Depth 2 then aggregates these per-epoch contributions over a sliding window of length $J$ and emphasizes their temporal variability (e.g., windowed variance). We maintain windowed statistics in a streaming manner (constant-time updates), so full trajectories need not be stored.
@@ E.3.3 Using Prediction Uncertainty with a Proxy (Dyn-Unc, DUAL)
Dyn-Unc (uncertainity_he2024_dynunc) measures prediction uncertainty via a sliding window of length $J$ over per-example target-class probabilities and averages the windowed uncertainty across proxy training.
DUAL (cho2025_DUAL) combines uncertainty with difficulty (window-mean prediction) and computes scores from an early stage of proxy training.
@@ Appendix F Comparison of CLD with Influence
The impact measured by $\mathtt{CLD}$ closely aligns with the “influence" of individual training samples on a model’s predictions that are measured by Training Data Attribution (TDA) methods. TDA methods have been widely employed for tasks such as debugging datasets, interpreting models, and optimizing training efficiency Koh2017; representer_yeh2018; FZ_infl_feldman2020.
The earliest TDA methods utilized Leave-One-Out (LOO) training, which involves retraining the model after removing specific data points and observing the changes in performance. While straightforward, LOO retraining is computationally prohibitive for modern deep learning models due to the need for multiple retraining cycles Koh2017. Recent TDA metrics, such as FZ-Influence ($\mathtt{Infl}$FZ_infl_feldman2020 and $\mathtt{Datamodels}$ datamodels_ilyas2022, have gained popularity owing to precomputed scores for widely-used datasets in computer vision. These methods, however, face scalability challenges.
A prominent alternative that arose was Influence Functions, which estimated the effect of downweighting individual samples using first-order (gradient) and second-order (Hessian) computations Koh2017; influence_fragile_basu2021 performed at the end of training. Methods like $\mathtt{RandSelect}$ Randselect_wojnowicz2016 and $\mathtt{Arnoldi}$ iterations Arnoldi_schioppa2022 improved computational efficiency by approximating the Hessian. Similarly, $\mathtt{TRAK}$ trak_2023park combined random projections, gradient-based methods, and ensembling to estimate the influence of training samples. However, these approaches often rely on strong assumptions, such as convergence to a unique optimal solution, which limits their applicability to neural networks. Additionally, Hessian computations introduce significant computational overhead. To address these challenges, unrolling-based methods that observe the learning process across training iterations have been proposed. These techniques approximate the impact of samples by differentiating through the optimization trajectory dataclensing_hara2019. Among these, $\mathtt{TracIn}$ tracin_pruthi2020 is a highly efficient method that estimates influence using gradients tracked throughout training. Its practical implementation, $\mathtt{TracInCP}$, uses intermediate checkpoints to alleviate computational burdens. While effective, unrolling methods require storing intermediate training states, leading to high storage and computational costs.
@@
We will now compare the impact measured by this metric ($\mathtt{CLD}_{\texttt{infl}}$) to the influence measured by TDA metrics, by utilizing the linear datamodeling score ($\mathtt{LDS}$) introduced by trak_2023park. $\mathtt{LDS}$ measures the correlation between group-level attribution scores ($\mathtt{CLD}_{\texttt{infl}}$ or influence) and their observed impact on model predictions when subsets of training data are used.
$\mathbf{LDS}$ definition For a query data point $z_{q}$, random subsets $\{\mathcal{S}_{j}\}_{j=1}^{C}$ are sampled from the training dataset, where each subset $\mathcal{S}_{j}$ contains $\lceil\alpha N\rceil$ points, with $\alpha\in(0,1)$ as the sampling ratio. Each subset $\mathcal{S}_{j}$ is used to retrain the model $R$ times with different initializations $\{\xi_{r}\}_{r=1}^{R}$ and training parameters $\lambda$, resulting in the model $\theta^{T}_{S_{j},\xi_{r}}$. This trained model is then used to compute a measurable quantity $f(\vec{z}_{q},\theta^{T}_{\mathcal{S}_{j},\xi_{r}})$. A group attribution score, $g_{\tau}(\vec{z}_{q},\mathcal{S}_{j},\mathcal{S})$, is calculated as $g_{\tau}(\vec{z}_{q},\mathcal{S}_{j},\mathcal{S})\coloneqq\sum_{\vec{z}\in\mathcal{S}_{j}}\tau(\vec{z}_{q},\vec{z},\mathcal{S}),$ where $\tau(\vec{z}_{q},\vec{z},\mathcal{S})$ is the attribution score for a training point $\vec{z}$ with respect to $\vec{z}_{q}$. The $\mathtt{LDS}$ is then obtained using Spearman’s rank spearman1904 correlation ($\rho_{s}$):
@@
Experimental Setup: We compared the $\mathtt{LDS}$ scores of $\mathtt{CLD}_{\texttt{infl}}$ against those of $\mathtt{TRAK}$, $\mathtt{Arnoldi}$, $\mathtt{TracIn}$, $\mathtt{Infl}$, and $\mathtt{Datamodels}$. Precomputed scores for $\mathtt{Infl}$ and $\mathtt{Datamodels}$ were used for the CIFAR-10 dataset cifar with ResNet-9 resnet, while 10 models were trained for $\mathtt{TRAK}$, $\mathtt{Arnoldi}$, $\mathtt{TracIn}$, and $\mathtt{CLD}_{\texttt{infl}}$. The evaluation employed $C=100$ random subsets, sampling ratios $\alpha$ ranging from $0.3$ to $\frac{N-1}{N}$, a query set of 200 samples, and $R=10$ seeds. The measurable quantity was the accuracy of query samples.
@@
Results and Observations: The results presented in Figure˜8(a) reveal that while the impact captured by $\mathtt{CLD}_{\texttt{infl}}$ is distinct from the influence measured by traditional TDA metrics, it aligns closely with methods such as $\mathtt{TracIn}$ and $\mathtt{TRAK}$ in terms of behavior while being resource-efficient. Notably, the performance gap between these computationally intensive methods and $\mathtt{CLD}_{\texttt{infl}}$ narrows as $\alpha$ increases. The drop in $\mathtt{LDS}$ scores at $\alpha=\frac{N-1}{N}$ is due to the stochastic nature of model retraining [^2] This observation is consistent with the findings of previous research revisitinglds_karthikeyan2021; source_bae2024..